OpenStudy (gabrielah96):

Find exact value of tan195 using half angle identities

OpenStudy (anonymous):

tan(195) = tan(390) = tan(30)

OpenStudy (anonymous):

\[\tan(30/2) = \sqrt{(\sin(30))/(1+\cos(30))}\]

OpenStudy (anonymous):

wait... there is NO square root on the previous post.

OpenStudy (anonymous):

tan(30/2) = (sin(30))/(1+cos(30))

OpenStudy (anonymous):

which simplifies to 2-sqrt(3)